OpenStudy (anonymous):

A right cylinder has a height of 7 inches and the base has a diameter of 6 inches . Determine the lateral area , in square inches , of the cylinder in terms of pie

OpenStudy (anonymous):

given h= 7 in and d = 6 in i.e. r = 6/2 = 3 in lateral area of right cylinder = \[2 \pi r h = 2 \times \pi \times 3 \times 7 = 42 \pi sq inches\] @C_gabriela

OpenStudy (anonymous):

are you in FLVS? I am